OpenStudy (cenaida):

solve | x\3 +1| = 6 A.{-7, 15} B.{-21, 15} C.{7, 15} D.{-15, 21}

jimthompson5910 (jim_thompson5910):

|x/3 + 1| = 6 x/3 + 1 = 6 or x/3 + 1 = -6 x/3 = 6-1 or x/3 = -6-1 x/3 = 5 or x/3 = -7 I'll let you finish

OpenStudy (cenaida):

-21, 15 so its b ?

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (cenaida):

Thanks!

jimthompson5910 (jim_thompson5910):

yw